OpenStudy (anonymous):

Need Algebra Help Please. Will Fan & Medal :) Factor each expression. 44. 6k^2 - 10kl + 4l^2 = 54. -2h^2 + 4h + 70 = 68. w^2 + 24w + 144 = 70. 25x^2 - 36 =

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

factor a 2 out of the first one, start with \[2(3k^2-5kl+2l^2)\] then, by some miracle, the second part factors as \[2(3k-2l)(k-l)\]

OpenStudy (anonymous):

that was the hardest one the last one is the easiest since \[a^2-b^2=(a+b)(a-b)\] you have \(a=5x,b=6\) plug them in

OpenStudy (anonymous):

\[ w^2 + 24w + 144\] is a perfect square, i.e it is the square of \(w+12\)